# Adaptive Human Trajectory Prediction via Latent Corridors

### v1

@@ 8.2 Ablation on Prompt Location/Size
We ablated the prompt location and method of combining with the input. For the input location, we experimented with combining the prompt with several parts of the input to $\mathcal{P}$, $[\mathbf{M}_{\tau-H:\tau},S]$: all of the input heatmaps $\mathbf{M}_{\tau-H:\tau}$, just the first input heatmap $\mathbf{M}_{0}$, just the segmentation map $S$, and all of the inputs. For method of combination, we experimented with element-wise summing and element-wise multiplication. We ran this ablation on the latent corridors-only approach to training on MOTSynth. Results can be seen in Table 2. While summing seems to lead to better performance than multiplying, and summing to the heatmaps seems to be more helpful than summing to the segmentation, generally, the prompts are effective at improving performance on a variety of input locations. This shows promise for training latent corridors in a variety of settings.
@@
Table 2: Ablation on 437 MOTSynth scenes in the ATP latent corridor adaptation configuration. We experiment with different locations and methods of combining the prompt with the input, and find that summing the prompt to all input heatmaps yields the best result, but most combinations result in an improvement on the scene-aware baseline (see Table 1).
Figure 8: Adaptation over time FDE. As in Fig. 5, the x-axis represents normalized adaptation time in person-seconds. The y-axis represents the FDE (lower is better). Results are normalized per-scene and averaged over models trained on 25 MOTSynth scenes (a) and 7 from MOT and WildTrack (b), with shaded area $\sigma/10$. For the FDE metric, our methods improve on the baselines increasingly with adaptation time. Latent corridors + per-scene finetuning has the best performance, as with FDE, and ATP via just finetuning or just latent corridor learning is still comparable. c) Comparison to baseline over many MOTSynth scenes for models trained with $8\%$ (point) and $80\%$ (arrowhead) human-second datasets for FDE. For many deployment scenes, FDE improves significantly more with our method than ADE improved, but still, the per-scene improvements are varied.
9 Additional Experimental Results on MOTSynth, MOT and WildTrack
@@
We showcase additional qualitative results on MOTSynth data in Fig. 9. We see that our approach is able to learn that in a scene with a large staircase, pedestrians mostly move towards the staircase, regardless of the direction of their observed history (top left and middle). We also see that our approach learns that pedestrians tend to stay on a walkway (top right). Finally, we see several examples of our approach having awareness of the 3D ground plane and predicting future trajectories that lie within the ground plane (bottom two rows).
Finally, we break down the numbers for the three ATP methods used on Wildtrack and MOT shown in Fig. 5a and Fig. 8a in Table 3. We see that on these real datasets, for both ADE and FDE, ATP via per-scene finetuning alone yields slightly better results than ATP via latent corridor adaption only. However, a combination of the two significantly outperforms either alone.
MOT Wildtrack
Method ADE FDE ADE FDE
ATP (Finetune) 41.8 86.9 31.7 63.5
ATP (LC) 43.0 89.2 32.9 65.9
ATP (LC + PSF) 37.4 74.3 27.4 54.7
Table 3: ATP via latent corridors, per-scene finetuning, or a combination thereof on MOT and Wildtrack. We see that on real data, adaptation via per-scene finetuning alone outperforms latent corridors alone. However, a combination of our latent corridor approach with per-scene finetuning yields significantly better results.
10 Additional Results on Webcam Data
We scraped two additional 5-minute videos from EarthCam as described in Sec. 5.1, one from Rick’s Cafe in Jamaica and one from Times Square in New York City. Qualitative results for these scenes are in Fig. 10. In the top row, at the cafe, there is a complicated path through an overlook that twists down stairs towards the water. The scene-aware baseline often predicts that people will jump over a ledge, whereas our method learns the boundaries of the paths that people follow and is able to predict that people will stay within those boundaries. In the bottom row, we see that the scene-aware baseline does not recognize that people on a billboard will stay within the billboard, whereas our method is able to recognize that.
Quantitative results for these scenes, as well as for the NoLA daytime and nighttime scenes, can be seen in Table 4. We see that across the four EarthCam scenes, ATP via per-scene finetuning alone works better than using latent corridor adaptation alone (by a narrow margin on both NoLA scenes, and by a signficant amount on the Rick’s Cafe scene), but a combination of the two is significantly better than any other adaptive or non-adaptive approach. Interestingly, for the Rick’s Cafe and Times Square scenes, a constant velocity baseline is better than our non-adaptive learned baselines, but our ATP approach outperforms all baselines.
Figure 10: Qualitative results on additional EarthCam scenes. (top) At a cafe in Jamaica with an overlook and stairs on the edge of the water, our model (orange) is able to correctly predict that people will stick to moving up and down the path created by the stairs, while the scene-aware baseline (purple) predicts people will walk directly into the water (left) or over the edge of the path (middle, right). (bottom) In Times Square, a billboard depicts human actors in motion. The scene-aware baseline assumes that these actors will move following their observed history, while our model correctly predicts that anyone in the area of the billboard will stay in the billboard.
Method Rick’s Cafe Times Square NoLA Daytime NoLA Nighttime
@@ 10 Additional Results on Webcam Data
Table 4: Results on four EarthCam scenes. Across all of these challenging in-the-wild scenarios, our ATP method using latent corridors and per-scene finetuning outperforms the baselines.

### v2

@@ 8.3 Ablation on Prompt Location/Size
We ablated the prompt location and method of combining with the input. For the input location, we experimented with combining the prompt with several parts of the input to $\mathcal{P}$, $[\mathbf{M}_{\tau-H:\tau},S]$: all of the input heatmaps $\mathbf{M}_{\tau-H:\tau}$, just the first input heatmap $\mathbf{M}_{0}$, just the segmentation map $S$, and all of the inputs. For method of combination, we experimented with element-wise summing and element-wise multiplication. We ran this ablation on the latent corridors-only approach to training on MOTSynth. Results can be seen in Table 3. While summing seems to lead to better performance than multiplying, and summing to the heatmaps seems to be more helpful than summing to the segmentation, generally, the prompts are effective at improving performance on a variety of input locations. This shows promise for training latent corridors to adapt a variety of architectures.
@@
Table 3: Ablation on 437 MOTSynth scenes in the ATP latent corridor adaptation configuration. We experiment with different locations and methods of combining the prompt with the input, and find that summing the prompt to all input heatmaps yields the best result, but most combinations result in an improvement on the scene-aware baseline (see main text Table 1).
Figure 8: Adaptation over time FDE. As in main text Fig. 5, the x-axis represents normalized adaptation time in human-seconds. The y-axis represents the FDE (lower is better). Results are normalized per-scene and averaged over models trained on 25 MOTSynth scenes (a) and 7 from MOT and WildTrack (b), with shaded area $\sigma/10$. For the FDE metric, our methods improve on the baselines increasingly with adaptation time. Latent corridors + per-scene finetuning has the best performance, as with FDE, and ATP via just finetuning or just latent corridor learning is still comparable. c) Comparison to baseline over many MOTSynth scenes for models trained with $8\%$ (point) and $80\%$ (arrowhead) human-second datasets for FDE. For many deployment scenes, FDE improves significantly more with our method than ADE improved, but still, the per-scene improvements are varied.
8.4 Segmentation Classes
We condense Mask2Former’s 132 classes into 12 classes that are meaningful for outdoor pedestrians: person, bicycle, car, motorcycle, large vehicle, traffic light, stop sign, bench, stairs, road/ground, building/wall, other.
9 Additional Experimental Results
9.1 Results on MOTSynth, WildTrack, MOT, and EarthCam
@@
We showcase additional qualitative results on MOTSynth data in Fig. 9. We see that our approach is able to learn that in a nighttime scene with a large staircase, pedestrians mostly move towards the staircase, regardless of the direction of their observed history (top left and middle). We also see that our approach learns that pedestrians tend to stay on a walkway (top right). Finally, we see several examples of our approach having awareness of the 3D ground plane and predicting future trajectories that lie within the ground plane (bottom two rows).
Quantitative results for each of the EarthCam scenes can be seen in Table 4. We see that across the four EarthCam scenes, ATP via per-scene finetuning alone works better than using latent corridor adaptation alone (by a narrow margin on both NoLA scenes, and by a significant amount on the Rick’s Cafe scene), but a combination of the two is significantly better than any other adaptive or non-adaptive approach. Interestingly, for the Rick’s Cafe and Times Square scenes, a constant velocity baseline is better than our non-adaptive learned baselines, but our ATP approach outperforms all baselines.
Method Rick’s Cafe Times Square NoLA (Day) NoLA (Night)
@@
Table 4: Results on four EarthCam scenes. Across all of these challenging in-the-wild scenarios, our ATP method using latent corridors and per-scene finetuning outperforms the baselines and other ATP methods.
9.2 Results on ETH/UCY
The main text focused on challenging datasets with non-top-down camera viewpoints, as compared to birds-eye-view datasets popular in earlier works such as SDD and ETH/UCY. Here, we run our approach in this setting by evaluating on ETH/UCY. For each scene in ETH/UCY (ETH, HOTEL, UNIV, ZARA1, ZARA2), we construct an 80/20 train-test split and evaluate as described in main text section 6.2. Results in Table 5 are in pixels and we compute ADE/FDE on a single predicted future. We see a $4.6\%$ improvement in ADE and $2.4\%$ in FDE using LC over per-scene finetuning.
Method ADE FDE
Scene Aware (YNet-Ours) 32.3 69.1
ATP (Per-Scene Finetune) 22.8 46.6
ATP (LC + Per-Scene Finetune) 21.8 45.5
Table 5: Results on ETH/UCY. ATP using latent corridors and per-scene finetuning outperforms ATP using fine-tuning alone, and both approaches successfully adapt over the scene-aware baseline.
10 Choice of Baselines
Our key scene-aware baseline is YNet, which outperforms or is comparable to other methods that utilize scene priors and trajectory histories such as [30, 29, 53, 31]. Since we used a simplified version of the YNet architecture for our scene-aware baseline, we have further benchmarked against the original YNet using the codebase training configuration using MOTSynth and the Stanford Drone Dataset [36]. We see in Table 6 that the YNet-Ours outperforms the original YNet in the unimodal setting.
SDD MOTSynth
Method ADE FDE ADE FDE
YNet [27] 24.9 49.9 54.4 112.3
YNet-Ours 17.3 33.6 47.3 96.5
Table 6: The original YNet model compared to our unimodal-only implementation (YNet-Ours). We see that on both SDD and MOTSynth, YNet-Ours outperforms YNet, and is therefore a stronger baseline.